# Better Instruction-Following Through Minimum Bayes Risk

### v3

@@ 1 Introduction
From our MBR inference experiments, we find that MBR decoding with LLM judge Prometheus-2-7B (Kim et al., 2024) improves performance by +3.6% on AlpacaEval and +0.28 on MT-Bench on average across five LLMs relative to greedy decoding. Notably, Llama2-7b with Prometheus MBR decoding outperforms Llama2-13b with greedy decoding on MT-Bench, while Prometheus MBR decoding with Llama2-13b outperforms Llama2-70b with greedy decoding on AlpacaEval 2.0. Gains persist even for large 70B models, demonstrating that small LLM judges can supervise larger LLMs through MBR decoding. We also compare MBR decoding against other methods that use LLM judges for supervision. We show that Prometheus MBR decoding is far more effective than MBR decoding with word match-based metrics (e.g. ROUGE) or semantic similarity-based metrics (e.g. BERTScore). Comparing MBR to BoN decoding, we find that MBR decoding consistently outperforms BoN decoding across multiple LLMs and LLM judges, and that the gains from MBR decoding increase as the supervising LLM judge increases in size and ability.
From our MBR distillation experiments, we find that self-training with Direct Preference Optimisation (DPO) (Rafailov et al., 2024) on preference pairs selected using MBR decoding (Yang et al., 2024) with Prometheus-2-7B substantially improves greedy decoding performance. For instance, MBR self-trained Llama2-13b improves by +7.1% on AlpacaEval 2.0 and +0.90 on MT-Bench relative to its baseline SFT counterpart when evaluated using only greedy decoding, far surpassing the corresponding gains from BoN self-training. We also find that MBR self-trained models evaluated with greedy decoding generally match and sometimes exceed the performance of their base models evaluated with MBR decoding, thereby demonstrating that MBR distillation is an effective way of mitigating the inference-time costs of MBR decoding while retaining improved performance.
@@ 2 Background
During inference, outputs are typically obtained either using maximum a-posteriori-based decoding methods that attempt to maximise probability, such as greedy decoding ($y_{i}=\operatorname*{arg\,max}_{y_{i}}p(y_{i}|y_{<i},x)$) and beam search (Graves, 2012), or by tokenwise sampling from the distribution ($y_{i}\sim p(y_{i}|y_{<i},x)$). Both approaches rely on the model’s distribution as indicative of output quality.
@@
where the expectation is approximated as a Monte Carlo sum using model samples $y^{(1)},\ldots,y^{(N_{\mathrm{cand}})}\sim p(y|x)$. In practice, this amounts to computing the utility of each candidate in $\mathcal{H}_{\mathrm{hyp}}$ using all other candidates as (pseudo-)references, and then selecting the candidate with the highest average utility as the final output. [^2] The expectation can also be computed over a separate set of model outputs known as the evidence set (Eikema & Aziz, 2020; Bertsch et al., 2023). We do not explore this setting in our work. The MBR-decoded output can therefore be interpreted as being the candidate with the highest “consensus” utility as measured by the utility metric, as it achieves the highest average utility when evaluated against all other candidate outputs. It is therefore crucial to choose a reference-based metric that is a good proxy for human preferences as our utility function, as this ensures that a “high-consensus” output corresponds to a “high-quality” output.
@@ 3.1.1 Models and Generation Parameters
We use the chat and instruct variants of the Llama2 (Touvron et al., 2023b) and Llama3 (Dubey et al., 2024) models in this experiment. All models have undergone prior SFT and demonstrate strong instruction-following and conversation abilities. We generate $N_{\mathrm{cand}}=30$ candidates using temperature sampling with $t=0.3$ for all MBR decoding experiments unless otherwise specified.
@@ 3.1.3 Baselines
In addition to greedy decoding and beam search (BS) with $k=10$ beams, we also experiment with longest decoding, where we select the longest candidate from the hypothesis set (as measured in characters) as the final output, and best-of-N (BoN) decoding. We generate $N_{\mathrm{cand}}=30$ candidates using temperature sampling with $t=0.3$ for both longest and BoN decoding. See Appendices A.2 and A.3 for comparison to additional baselines.
@@ 3.2 Experimental Results
Prometheus MBR decoding provides significant and consistent gains The gains associated with Prometheus MBR decoding are significantly larger than those associated with other utility metrics, yielding an average improvement of +3.6% on AlpacaEval 2.0 and +0.28 on MT-Bench. For comparison, the performance gap between Llama2-7b and Llama2-13b with greedy decoding is +4.6% on AlpacaEval 2.0 and +0.18 on MT-Bench, while the corresponding gap between Llama2-13b and Llama2-70b is +3.8% and +0.60. Notably, Llama2-7b with Prometheus MBR decoding outperforms Llama2-13b with greedy decoding on MT-Bench, while Prometheus MBR decoding with Llama2-13b outperforms Llama2-70b - a model over five times bigger - with greedy decoding on AlpacaEval 2.0. We also find that Prometheus MBR decoding yields larger gains than Prometheus BoN decoding; we explore this further in Section 3.3.1.
We also highlight that the performance gains associated with Prometheus MBR decoding are significant across models of all sizes, even for much larger models such as Llama3-70b. This scaling property suggests that small judge models can still be used to supervise much larger models.
@@ 3.3 Analysis of Prometheus MBR decoding
Figure 2: AlpacaEval 2.0 win rates (%) for Llama2-70b with varying hypothesis set size $N_{\mathrm{cand}}$ (left) and generation temperature $t$ (right) values for Prometheus MBR and BoN decoding. Performance for both methods initially increases with $N_{\mathrm{cand}}$ and plateaus at around $N_{\mathrm{cand}}=20$. Performance also initially increases with $t$, but drops rapidly after $t=1.0$.
Prometheus MBR decoding performance vs. $t$ and $N_{\mathrm{cand}}$  We plot AlpacaEval 2.0 win rates as a function of $N_{\mathrm{cand}}$ and $t$ in Figure 2 for Llama2-70b with Prometheus MBR and BoN decoding. We find that performance initially increases with $N_{\mathrm{cand}}$ but plateaus at around $N_{\mathrm{cand}}=20$, suggesting that expanding the size of the hypothesis set beyond this yields little benefit. Performance also initially increases with $t$, highlighting the benefits of increased candidate diversity, although it rapidly degrades at high temperatures as the individual candidate outputs decline in quality.
@@ 3.3.1 Further Comparisons with Best-of-N Decoding
Llama3-70b-Instruct-BoN 5.77 6.16 6.57 7.39 8.35 0.06
Llama3-70b-Instruct-MBR 6.22 6.43 6.94 7.87 8.52 0.41
@@
As BoN decoding can also leverage LLM judges as a utility metric, we conduct additional experiments to compare its performance against MBR decoding. We compare BoN and MBR decoding for five different LLM judges on MT-Bench and report the results in Table 3. In addition to Prometheus-2-7B, we also evaluate its larger sibling Prometheus-2-8x7B, as well as JudgeLM-7b and JudgeLM-33b (Zhu et al., 2023c), which are two judge models finetuned from LLaMA models (Touvron et al., 2023a). We also assess Llama3-70b-Instruct as a zero-shot judge for MBR decoding (see Appendix E for our prompts). All of our chosen judges can act as both reference-free and reference-based judges, allowing us to compare MBR and BoN decoding on a level playing field. [^3] Because sequence-classifier reward models (Stiennon et al., 2022) do not support reference-based evaluation, it is not possible to fairly compare BoN decoding with these methods to MBR. We therefore do not discuss this in the main text and report our findings in Appendix F instead.
We find that MBR decoding consistently outperforms BoN decoding across all selected judge models. This difference is especially large for the JudgeLM models and for Llama3-70b-Instruct, where BoN fails to significantly improve on greedy decoding. One explanation for this discrepancy is that our LLM judges are insufficiently good at reference-free evaluation for BoN decoding to be effective. This idea is supported by prior studies comparing reference-free and reference-based evaluation, which consistently show that reference-free methods tend to underperform, even when using strong judge models like GPT-4 (Ye et al., 2024; Kim et al., 2024; Zheng et al., 2023). Another explanation is that MBR decoding provides a smoothing effect that arises from our use of expected utility in place of utility point estimates for output selection, tying back to our hypothesis that selecting “high-consensus” outputs yields significant benefit. This averaging process reduces the impact of individual mistakes made by the imperfect LLM judge, thereby providing for a more stable and reliable measure of quality. We leave further exploration of these ideas to future work.
Notably, in Table 3, MBR performance improves by scaling the size of the LLM judge, with Prometheus-2-8x7B outperforming Prometheus-2-7B and JudgeLM-33b outperforming JudgeLM-7b. This suggests that improving the LLM judge utility metric directly improves MBR decoding performance and that MBR decoding will benefit as newer and better LLM judges are developed.
@@ 4 MBR Distillation
Our results so far demonstrate the potential of MBR decoding to significantly improve the test-time performance of instruction-following models, but this comes at the cost of substantial inference-time compute costs due to the linear cost for generating $N_{\mathrm{cand}}$ candidate outputs, and the quadratic cost for computing utility across these candidates. To mitigate this, we explore distilling MBR-decoded outputs back into the model itself and aim to obtain MBR decoding-level (or better) performance without needing to perform MBR decoding at test time.
@@ 4.1.1 Training an SFT Model
We start by performing SFT on the base Llama2-7b and Llama2-13b models. This is necessary to instil instruction-following behaviour in the models so that they can be used to generate instruction-following self-training data. We choose not to use the official chat variants of these models as we wish to retain full control over the training procedure and avoid inheriting any biases introduced through prior finetuning and alignment.
We use 3000 randomly-drawn samples from UltraChat-200k (Ding et al., 2023) for SFT. UltraChat is a highly diverse conversational instruction-following dataset created using GPT-3.5-Turbo. Each sample consists of multi-turn prompts and responses, although we only take the first turn of each sample in order to simplify experimentation. We designate our SFT models as sft.
@@ 4.1.2 Iterative DPO on MBR-Decoded Outputs
Having obtained SFT models, we now conduct DPO to improve the models on their own MBR-decoded outputs, an approach first proposed by Yang et al. (2024) for improving machine translation. We start by randomly drawing a further 3000 prompts from UltraChat (excluding the samples that have already been selected for SFT). Next, we generate $N_{\mathrm{cand}}=12$ candidate outputs from our sft models using these prompts. Following Yang et al. (2024), we then score the candidate outputs using a utility metric and form preference pairs from the highest-scoring and lowest-scoring outputs. This preference pair dataset is then used for DPO on the sft models, yielding dpo-MBR-1. We extend upon Yang et al. (2024) by iteratively repeating this process twice more, each time using the latest dpo models as the base model paired with a fresh set of 3000 prompts, yielding the models dpo-MBR-2 and dpo-MBR-3. See Appendix I for a summary of our SFT and DPO hyperparameters and Appendix G.4 for experimental results from using another preference pair selection strategy.
@@ 4.1.3 Utility Metrics and Evaluation
We use Prometheus-2-7B as our utility metric, as this yielded the most promising results in our earlier experiments (Section 3.2), although we also try MBR self-training with ROUGE as the utility metric (Appendix G.3). We compare our dpo models with greedy decoding against the sft models with greedy decoding, beam search and MBR decoding. For MBR decoding, we use $N_{\mathrm{cand}}=30$ and $t=0.3$ with Prometheus-2-7B as the utility metric. We also baseline against models trained with SFT on 12000 UltraChat samples that we call sft-full. Finally, we experiment with BoN self-training, again using Prometheus-2-7B as the utility metric and following the same procedure as for MBR self-training, which yields the models dpo-BoN-1, dpo-BoN-2 and dpo-BoN-3.
@@ 5 Related Work
Inference-time algorithms Many inference-time algorithms generate multiple candidate outputs and select a final output based on external criteria. In addition to MBR and BoN decoding, examples of such algorithms include Self-Consistency (Wang et al., 2023), which selects the most self-consistent answers through marginalisation of chain-of-thought reasoning paths and Universal Self-Consistency (USC) (Chen et al., 2023), where the LLM is used to self-select the most consistent chain-of-thought reasoning path from amongst many reasoning paths. Kuhn et al. (2023) propose an MBR-esque algorithm that uses dense embedders and clustering to measure semantic uncertainty. Other inference-time algorithms prompt the LLM to perform additional inference steps in a structured manner. Examples include Tree-of-Thoughts (Yao et al., 2023) and Graph-of-Thoughts (Besta et al., 2024), as well as recursive improvement strategies such as Self-Refine (Madaan et al., 2023) and Reflexion (Shinn et al., 2023).
Self-training Self-training is a promising avenue for model improvement as it enables training without external labelled data. Gulcehre et al. (2023) introduce an algorithm that iteratively generates samples from a policy and then updates the policy using offline RL. Yuan et al. (2024b) train models to score their own outputs, and then use these scores to create preference datasets which they distil back into the model using DPO. Huang et al. (2022) train models on their own highest-confidence outputs as determined by majority voting. Self-training on MBR-decoded outputs has also been explored for machine translation. Finkelstein et al. (2024) train models with SFT on their own MBR and quality estimation outputs for machine translation and demonstrate that this yields improvements over baseline models. Wang et al. (2024a) use MBR to generate targets for sequence-level distillation, again for machine translation. Yang et al. (2024) are the first to use DPO to upweight the model’s own MBR generations, allowing them to recover much of their original MBR performances on machine translation using only greedy decoding.

### v4

@@ 1 Introduction
From our MBR inference experiments, we find that MBR decoding with LLM judge Prometheus-2-7B (Kim et al., 2024) improves performance by +3.6% on AlpacaEval and +0.28 on MT-Bench on average across five LLMs relative to greedy decoding. Notably, Llama-2-7b with Prometheus MBR decoding outperforms Llama-2-13b with greedy decoding on MT-Bench, while Prometheus MBR decoding with Llama-2-13b outperforms Llama-2-70b with greedy decoding on AlpacaEval 2.0. Gains persist even for large 70B models, demonstrating that small LLM judges can supervise larger LLMs through MBR decoding. We also compare MBR decoding against other methods that use LLM judges for supervision. We show that Prometheus MBR decoding is far more effective than MBR decoding with word match-based metrics (e.g. ROUGE) or semantic similarity-based metrics (e.g. BERTScore). Comparing MBR to BoN decoding, we find that MBR decoding consistently outperforms BoN decoding across multiple LLMs and LLM judges, and that the gains from MBR decoding increase as the supervising LLM judge increases in size and ability.
From our MBR distillation experiments, we find that self-training with Direct Preference Optimisation (DPO) (Rafailov et al., 2024) on preference pairs selected using MBR decoding (Yang et al., 2024) with Prometheus-2-7B substantially improves greedy decoding performance. For instance, MBR self-trained Llama-2-13b improves by +7.1% on AlpacaEval 2.0 and +0.90 on MT-Bench relative to its baseline SFT counterpart when evaluated using only greedy decoding, far surpassing the corresponding gains from BoN self-training. We also find that MBR self-trained models evaluated with greedy decoding generally match and sometimes exceed the performance of their base models evaluated with MBR decoding, thereby demonstrating that MBR distillation is an effective way of mitigating the inference-time costs of MBR decoding while retaining improved performance.
@@ 2 Background
During inference, outputs are typically obtained either using maximum a-posteriori-based decoding methods that attempt to maximise probability, such as greedy decoding ($y_{i}=\operatorname*{arg\,max}_{y_{i}}p(y_{i}|y_{<i},x)$) and beam search (Graves, 2012), or by tokenwise sampling from the distribution ($y_{i}\sim p(y_{i}|y_{<i},x)$). Both rely on the model’s distribution as indicative of output quality.
@@
where the expectation is approximated as a Monte Carlo sum using model samples $y^{(1)},\ldots,y^{(N_{\mathrm{cand}})}\sim p(y|x)$. In practice, this amounts to computing the utility of each candidate in $\mathcal{H}_{\mathrm{hyp}}$ using all other candidates as (pseudo-)references, and then selecting the candidate with the highest average utility as the final output [^2] The expectation can also be computed over a separate set of model outputs known as the evidence set (Eikema & Aziz, 2020; Bertsch et al., 2023). We do not explore this setting in our work. - see Appendix I.1 for an algorithmic description of MBR decoding. The MBR-decoded output can therefore be interpreted as being the candidate with the highest “consensus” utility as measured by the utility metric, as it achieves the highest average utility when evaluated against all other candidate outputs. It is therefore crucial to choose a reference-based metric that is a good proxy for human preferences as our utility function, as this ensures that a “high-consensus” output corresponds to a “high-quality” output.
@@ 3.1.1 Models and Generation Parameters
We use the chat and instruct variants of the Llama-2 (Touvron et al., 2023b) and Llama-3 (Dubey et al., 2024) models in this experiment. All models have undergone prior SFT and demonstrate strong instruction-following and conversation abilities. We generate $N_{\mathrm{cand}}=30$ candidates using temperature sampling with $t=0.3$ for all MBR decoding experiments unless otherwise specified.
@@ 3.1.3 Baselines
In addition to greedy decoding and beam search (BS) with $k=10$ beams, we also experiment with longest decoding, where we select the longest candidate from the hypothesis set (as measured in characters) as the final output, and best-of-N (BoN) decoding. We generate $N_{\mathrm{cand}}=30$ candidates using temperature sampling with $t=0.3$ for both longest and BoN decoding. See Appendices A.2 and A.3 for comparisons with additional baselines.
@@ 3.2 Experimental Results
Prometheus MBR decoding provides significant and consistent gains The gains associated with Prometheus MBR decoding are significantly larger than those associated with other utility metrics, yielding an average improvement of +3.6% on AlpacaEval 2.0 and +0.28 on MT-Bench. For comparison, the performance gap between Llama-2-7b and Llama-2-13b with greedy decoding is +4.6% on AlpacaEval 2.0 and +0.18 on MT-Bench, while the corresponding gap between Llama-2-13b and Llama-2-70b is +3.8% and +0.60. Notably, Llama-2-7b with Prometheus MBR decoding outperforms Llama-2-13b with greedy decoding on MT-Bench, while Prometheus MBR decoding with Llama-2-13b outperforms Llama-2-70b - a model over five times bigger - with greedy decoding on AlpacaEval 2.0. We also find that Prometheus MBR decoding yields larger gains than Prometheus BoN decoding; we explore this further in Section 3.3.1.
We also highlight that the performance gains associated with Prometheus MBR decoding are significant across models of all sizes, even for much larger models such as Llama-3-70b. This scaling property suggests that small judge models can still be used to supervise much larger models.
@@ 3.3 Analysis of Prometheus MBR decoding
Figure 2: AlpacaEval 2.0 win rates (%) for Llama-2-70b with varying hypothesis set size $N_{\mathrm{cand}}$ (left) and generation temperature $t$ (right) values for Prometheus MBR and BoN decoding. Performance for both methods initially increases with $N_{\mathrm{cand}}$ and plateaus at around $N_{\mathrm{cand}}=20$. Performance also initially increases with $t$, but drops rapidly after $t=1.0$.
Prometheus MBR decoding performance vs. $t$ and $N_{\mathrm{cand}}$  We plot AlpacaEval 2.0 win rates as a function of $N_{\mathrm{cand}}$ and $t$ in Figure 2 for Llama-2-70b with Prometheus MBR and BoN decoding. We find that performance initially increases with $N_{\mathrm{cand}}$ but plateaus at around $N_{\mathrm{cand}}=20$, suggesting that expanding the size of the hypothesis set beyond this yields little benefit. Performance also initially increases with $t$, highlighting the benefits of increased candidate diversity, although it rapidly degrades at high temperatures as the individual candidate outputs decline in quality.
@@ 3.3.1 Further Comparisons with Best-of-N Decoding
Llama-3-8b-Instruct-BoN 5.83 6.05 6.61 7.60 8.38 0.10
Llama-3-8b-Instruct-MBR 5.96 6.28 6.84 7.80 8.47 0.28
Llama-3-70b-Instruct-BoN 5.77 6.16 6.57 7.39 8.35 0.06
Llama-3-70b-Instruct-MBR 6.22 6.43 6.94 7.87 8.52 0.41
@@
As BoN decoding can also leverage LLM judges as a utility metric, we conduct additional experiments to compare its performance against MBR decoding. We compare BoN and MBR decoding for five different LLM judges on MT-Bench and report the results in Table 3. In addition to Prometheus-2-7B, we also evaluate its larger sibling Prometheus-2-8x7B, as well as JudgeLM-7b and JudgeLM-33b (Zhu et al., 2023c), which are two judge models finetuned from LLaMA models (Touvron et al., 2023a). We also assess Llama-3-8b-Instruct and Llama-3-70b-Instruct as zero-shot judges for MBR decoding (see Appendix E for our prompts). All chosen judges can act as both reference-free and reference-based judges, allowing us to compare MBR and BoN decoding fairly. [^3] Because sequence-classifier reward models (Stiennon et al., 2022) do not support reference-based evaluation, it is not possible to fairly compare BoN decoding with these methods to MBR. We therefore do not discuss this in the main text and report our findings in Appendix F instead.
We find that MBR decoding consistently outperforms BoN decoding across all selected judge models. This difference is especially large for the JudgeLM models and for Llama-3-70b-Instruct, where BoN fails to significantly improve on greedy decoding. One explanation for this discrepancy is that our LLM judges are insufficiently good at reference-free evaluation for BoN decoding to be effective. This idea is supported by prior studies comparing reference-free and reference-based evaluation, which consistently show that reference-free methods tend to underperform, even when using strong judge models like GPT-4 (Ye et al., 2024; Kim et al., 2024; Zheng et al., 2023). Another explanation is that MBR decoding provides a smoothing effect that arises from our use of expected utility in place of utility point estimates for output selection, tying back to our hypothesis that selecting “high-consensus” outputs yields significant benefit. This averaging process reduces the impact of individual mistakes made by the imperfect LLM judge, thereby providing for a more stable and reliable measure of quality. We leave further exploration of these ideas to future work.
Notably, in Table 3, MBR performance improves by scaling the size of the LLM judge, with Prometheus-2-8x7B outperforming Prometheus-2-7B, JudgeLM-33b outperforming JudgeLM-7b, and Llama-3-70b-Instruct outperforming Llama-3-8b-Instruct. This suggests that improving the LLM judge utility metric directly improves MBR decoding performance and that MBR decoding will benefit as newer and better LLM judges are developed.
@@ 4 MBR Distillation
Our results so far demonstrate the potential of MBR decoding to significantly improve the test-time performance of instruction-following models, but this comes at the cost of substantial inference-time compute costs due to the linear cost for generating $N_{\mathrm{cand}}$ candidate outputs and the quadratic cost for computing utility across these candidates. To mitigate this, we explore distilling MBR-decoded outputs back into the model itself and aim to obtain MBR decoding-level (or better) performance without needing to perform MBR decoding at test time.
@@ 4.1.1 Training an SFT Model
We start by performing SFT on the base Llama-2-7b and Llama-2-13b models. This is necessary to instil instruction-following behaviour in the models so that they can be used to generate instruction-following self-training data. We choose not to use the official chat variants of these models as we wish to retain control over the training procedure and avoid inheriting any biases introduced through prior finetuning and alignment. We use 3000 random samples from UltraChat (Ding et al., 2023) for SFT. UltraChat is a diverse conversational instruction-following dataset created using GPT-3.5-Turbo. Each sample consists of multi-turn prompts and responses, although we only take the first turn of each sample in order to simplify experimentation. We designate our SFT models as sft.
@@ 4.1.2 Iterative DPO on MBR-Decoded Outputs
Having obtained SFT models, we now conduct DPO to improve the models on their own MBR-decoded outputs, an approach first proposed by Yang et al. (2024) for improving machine translation. We start by randomly drawing a further 3000 prompts from UltraChat (excluding the samples that have already been selected for SFT). Next, we generate $N_{\mathrm{cand}}=12$ candidate outputs from our sft models using these prompts. We use a smaller $N_{\mathrm{cand}}$ than for MBR inference to balance performance and compute cost, as we know from Figure 2 that using an $N_{\mathrm{cand}}$ value above 10 already yields significant gains. Following Yang et al. (2024), we then score the candidate outputs using a utility metric and form preference pairs from the highest-scoring and lowest-scoring outputs. This preference pair dataset is then used for DPO on the sft models, yielding dpo-MBR-1. We extend upon Yang et al. (2024) by iteratively repeating this process twice more, each time using the latest dpo models as the base model paired with a fresh set of 3000 prompts, yielding the models dpo-MBR-2 and dpo-MBR-3. See Appendix K for a summary of our SFT and DPO hyperparameters, Appendix G.5 for experimental results from using another preference pair selection strategy, and Appendix I.2 for mathematical and algorithmic overviews of MBR distillation.
@@ 4.1.3 Utility Metrics and Evaluation
We use Prometheus-2-7B as our utility metric, although we also try MBR self-training with ROUGE and Llama-3-8b-Instruct as the utility metrics (Appendix G.3 and G.4). We compare our dpo models with greedy decoding against the sft models with greedy decoding, beam search and MBR decoding. For MBR decoding, we use $N_{\mathrm{cand}}=30$ and $t=0.3$ with Prometheus-2-7B as the utility metric. We also baseline against models trained with SFT on 12000 UltraChat samples (sft-full). Finally, we experiment with BoN self-training, again using Prometheus-2-7B as the utility metric and following the same procedure as for MBR self-training, which yields the models dpo-BoN-1, dpo-BoN-2 and dpo-BoN-3.
@@ 5 Related Work
Inference-time algorithms Many inference-time algorithms generate candidate outputs and select a final output based on external criteria. In addition to MBR and BoN decoding, examples include Self-Consistency (Wang et al., 2023), which selects the most self-consistent answers through marginalisation of chain-of-thought reasoning paths and Universal Self-Consistency (USC) (Chen et al., 2023), where the LLM is used to self-select consistent chain-of-thought reasoning paths from amongst many reasoning paths. Kuhn et al. (2023) propose an MBR-esque algorithm that uses dense embedders and clustering to measure semantic uncertainty. Other inference-time algorithms prompt the LLM to perform additional inference steps in a structured manner. Examples include Tree-of-Thoughts (Yao et al., 2023) and Graph-of-Thoughts (Besta et al., 2024), as well as recursive improvement strategies such as Self-Refine (Madaan et al., 2023) and Reflexion (Shinn et al., 2023).
Self-training Self-training is a promising avenue for model improvement as it enables training without labelled data. Gulcehre et al. (2023) introduce an algorithm that generates samples from a policy and then updates the policy using offline RL. Yuan et al. (2024b) train models to score their own outputs, and then use these scores to create preference datasets which for distillation. Huang et al. (2022) train models on their own highest-confidence outputs as determined by majority voting. Self-training on MBR-decoded outputs has also been explored for machine translation. Finkelstein et al. (2024) train models with SFT on their own MBR and quality estimation outputs for machine translation and demonstrate that this yields improvements over baseline models. Wang et al. (2024a) use MBR to generate targets for sequence-level distillation, again for machine translation. Yang et al. (2024) are the first to use DPO to upweight the model’s own MBR generations, allowing them to recover much of their original MBR performances on translation using only greedy decoding.